# Simplified and Generalized Masked Diffusion for Discrete Data

### v1

@@ F.1 The CTMC point of view
Proof    Consider the backward transition from time $t+\tau$ to $t$. For $j\neq k$, Bayes’ rule yields
@@
We use the shorthand $R_{\theta}(t)_{kj}$ to denote the approximate backward transition rate from the state $k$ to $j$ obtained by substituting our prediction model $\mu_{\theta}(k)$ for $x_{0}$ in $R^{x_{0}}(t)_{kj}$. Then, the continuous-time objective (4) can be equivalently expressed as
@@
F.2 Differences from Campbell et al. [29]
@@
Recall that $R_{\theta}(t)_{jk}$ is computed as substituting our neural network prediction model $\mu_{\theta}(j)$ for $x_{0}$ in $R^{x_{0}}(t)_{jk}$. Therefore, the summation together with $R_{\theta}(t)_{kk}$ requires $m$ evaluations of $\mu_{\theta}(\cdot)$. This is prohibitive since the neural network model is usually expensive. To resolve this issue, Campbell et al. [29] proposed to rewrite the sum as
@@ F.3 Score parameterization
Note that only the result under the case $k=m$ is needed. This is because when $x_{t}$ is unmasked, at any time between 00 and $t$, the state must stay unchanged and remain $x_{0}$. As a result, $\mathrm{KL}({q(x_{t-\Delta t}|x_{t},x_{0})}\|{p_{\theta}(x_{t-\Delta t}|x_{t})})=0$ for $x_{t}\neq m$. From (17), we know $Q(t)_{jk}=\beta(t)(\delta_{mk}-\delta_{jk})$. Combining (26) and (9), we get
@@
we conclude that this recovers the objective in (4). Interestingly, in Lou et al. [32] the score parameterization is not constrained to satisfy (28). That means the learned backward model might be incompatible with the forward process.
It is easy to prove Proposition 1 from Equation 26.
@@
F.4 Differences from Sundae
Sundae [59, 60] uses a linear noise schedule and uniformly corrupts data with random tokens in the vocab (Uniform style). Additionally it uses a second loss term from cross entropy between un-corrupted data and 1-step unrolled model prediction. Similar ideas have been proposed in [61].
Appendix G Details for state-dependent rates
G.1 Derivations and time continuous limit
All derivations in this section assume that $x_{t}$ is a single token, while for $N$ tokens the masked diffusion with state-dependent rates factorises across the $N$ tokens. Learning from data of $N$ tokens using variational inference is discussed in Section G.2.
Given the forward transition $q(x_{t}|x_{s})$ and marginal $q(x_{s}|x_{0})$ derived in main text (Section 5) The reversal given $x_{0}$ is $q(x_{s}|x_{t},x_{0})=\mathrm{Cat}(x_{s};\bar{R}^{x_{0}}(t,s)^{\top}x_{t})$ for
@@ G.1 Derivations and time continuous limit
Then, combining (30) and (31) to write $q(x_{s}|x_{t}=m,x_{0})$ in an unified way yields the expression (13) in the main Section 5. In the second case, when $x_{t}=x_{0}\neq m$, $q(x_{s}|x_{t}\neq m,x_{0})$ from (29) simplifies dramatically and it becomes $q(x_{s}|x_{t}\neq m,x_{0})=x_{t}^{\top}x_{s}$ which is a point mass that sets $x_{s}=x_{t}$.
@@
G.2 Training and gradient estimation
The model is applied to data consisted of $N$ tokens where $x_{0}=(x_{0}^{1},\ldots,x_{0}^{(N)})$ and where each state in the masked diffusion is $x_{t}=(x_{t}^{1},\ldots,x_{t}^{(N)})$. The backward generated model has a factorizing transition conditional of the form $\prod_{n=1}^{N}p_{\theta}(x_{s}^{(n)}|x_{t})$ where $p_{\theta}(x_{s}^{(n)}|x_{t})=q(x_{s}^{(n)}|x_{t}^{(n)},\mu_{\theta}^{(n)}(x_{t},t))$ has a form that depends on whether $x_{t}^{(n)}=m$ or $x_{t}^{(n)}\neq m$. For the first case:
@@ G.2 Training and gradient estimation
where $\mu^{(n)}_{\theta}(x_{t},t)=\text{softmax}(f_{\theta}(x_{t}))$ is a $m+1$ dimensional probability vector modelled by a NN (where the final value is constrained to be zero since $\mu^{(n)}_{\theta}(x_{t},t)$ is a reconstruction of $x_{0}^{(n)}$ which cannot be mask, so in practice the NN classifier needs to have a softmax output only over the $m$ actual token classes). Crucially, note that the NN classifier receives as input the full state $x_{t}$ of all tokens, while additional time features to encode $t$ are also included. When $x_{t}^{(n)}\neq m$ the backward transition model is set to be $p_{\theta}(x_{s}|x_{t}^{(n)}\neq m,\{x_{t}^{(k)}\}_{k\neq n})=(x_{t}^{(n)})^{\top}x_{s}^{(n)}$ which matches precisely $q(x_{s}^{(n)}|x_{t}^{(n)}=m,x_{0}^{(n)})=(x_{t}^{(n)})^{\top}x_{s}^{(n)}$ from the forward process.
@@
Given that each $\alpha_{t,i}=1-t^{w_{i}}$, the backward model becomes
@@
$-\frac{1}{2t}\left(\nabla_{w}\log q(x_{t}^{1}|x_{0})-\nabla_{w}\log q(x_{t}^{1}|x_{0})\right)\left[f(x_{t}^{1},x_{0})-f(x_{t}^{2},x_{0})\right].$
@@
Appendix H Additional Results and Experimental Details
Figure 5: More unconditional samples from MD4 trained on ImageNet 64x64.
@@ Appendix H Additional Results and Experimental Details
We used the continuous-time ELBO and drew one sample of $t$ for each data to estimate the integral. To reduce the variance of training, we used the same antithetic sampling trick described in Kingma et al. [33] for continuous diffusion models. We used the linear masking schedule $\alpha_{t}=1-t$ and added a small shift $\epsilon=10^{-4}$ when $t$ is close to 00 and $1$ to ensure numerical stability. The shifted schedule is $\alpha_{t}=(1-2\epsilon)(1-t)+\epsilon$. The shift leads to a support mismatch between $q(x_{1}|x_{0})$ and the prior $p(x_{1})$, leading to an undefined KL divergence term. We explain in appendix E how to modify the prior distribution to allow small uniform probabilities in non-mask states to mitigate this problem. The shift leads to a non-zero reconstruction term and KL divergence term for the prior distribution but both are of negligible scale so we can safely ignore them when reporting the ELBO.
@@
H.2 OpenWebText
Additional unconditional generation from our MD4 Medium model is shown in Appendix I.
@@ H.2 OpenWebText
Figure 6: Sample quality evaluated by GPT-2 Large following Lou et al. [32]. We compare MD4 against GPT-2 (auto-regressive baseline) and SEDD (state of the art discrete diffusion model) in generating coherent sequences (1024 tokens). We investigate the effects of two orthogonal factors on sample quality: model size and decoding steps. The numbers for GPT-2 and SEDD are from Lou et al. [32].
H.3 Images
Additional unconditional generation from our MD4 model is presented in Figure 5.
We used the same linear masking schedule as in previous experiments in all reported results. We have observed that the cosine schedule leads to better sample quality so we used it instead for visualizing samples from the model.
We used the same U-Net plus self-attention architectures from the continuous diffusion model described in Kingma et al. [33] for CIFAR-10, except that we did not use Fourier feature inputs and added an additional input embedding layer with embedding size the same as the hidden dimension of the model. For ImageNet $64\times 64$, we reduced the number of residual blocks from 64 to 48 and add 12 additional diffusion transformer [64] blocks with 768 hidden dimension and 12 heads in the middle.
We used AdamW optimizer with learning rate 0.0002 and weight decay factor 0.01 for both datasets. The learning rate follows a cosine annealing after 100 warm up steps. The batch size is 128 for CIFAR-10 and 512 for ImageNet 64$\times$64. Our CIFAR-10 model is trained on 16 TPU-v5 lite for 24 hours. Our ImageNet-$64\times 64$ model is trained on 256 TPU-v4 for 2 days.
Appendix I Additional unconditional generation from MD4-M
I.1 MD4-M sample 1: 1024 tokens

### v2

@@ H.1 The CTMC point of view
Proof    Consider the reverse transition from time $t+\tau$ to $t$. For $j\neq k$, Bayes’ rule yields
@@
We use the shorthand $R_{\theta}(t)_{kj}$ to denote the approximate reverse transition rate from the state $k$ to $j$ obtained by substituting our prediction model $\mu_{\theta}(k)$ for $x_{0}$ in $R^{x_{0}}(t)_{kj}$. Then, the continuous-time objective (4) can be equivalently expressed as
@@
H.2 Differences from Campbell et al. [29]
@@
For single dimension case, the sum is analytically computable due to the sparse structure of $R_{\theta}(t)$—if $x_{t}=k$ is mask, the second term disappears; otherwise the only possible neighbor $j$ is a mask. However, for multidimensional data, $j$ will represent all $N-1$ neighbors in the forward process, i.e., the states we get from mask out a single unmasked dimension of $x_{t}=k$. Recall that $R_{\theta}(t)_{jk}$ is computed as substituting our neural network prediction model $\mu_{\theta}(j)$ for $x_{0}$ in $R^{x_{0}}(t)_{jk}$. Therefore, the summation together with $R_{\theta}(t)_{kk}$ requires $N$ evaluations of $\mu_{\theta}(\cdot)$. This is prohibitive since the neural network model is usually expensive. To resolve this issue, Campbell et al. [29] proposed to rewrite the sum as
@@ H.3 Score parameterization
Note that only the result under the case $k=m$ is needed. This is because when $x_{t}$ is unmasked, at any time between 00 and $t$, the state must stay unchanged and remain $x_{0}$. As a result, $\mathrm{KL}({q(x_{t-\Delta t}|x_{t},x_{0})}\|{p_{\theta}(x_{t-\Delta t}|x_{t})})=0$ for $x_{t}\neq m$. From (15), we know $Q(t)_{jk}=\beta(t)(\delta_{mk}-\delta_{jk})$. Combining (24) and (H.3), we get
@@
we conclude that this recovers the objective in (4). Interestingly, in Lou et al. [32] the score parameterization is not constrained to satisfy (26). That means the learned reverse model might be incompatible with the forward process.
Below, we prove Proposition 1 using the result from Eq. 24.
@@
H.4 Other related work.
MaskGIT [39].
MaskGIT is a diffusion-inspired iterative denoising model for discrete image tokens obtained through models such as VQ-VAE [71]. Training of MaskGIT follows the steps: (a) Sample $t\in[0,1]$. (b) Given a mask scheduling function $\gamma(t)$, sample $\gamma(t)N$ tokens to place masks. (c) For data $x_{0}$ of size $(m+1)\times N$ and the partially masked state $x_{t}$, minimize the negative log-likelihood
Our forward process satisfies $q_{t|0}(m|x_{0})=1-\alpha_{t}$. Therefore, when we set the mask scheduling function as $\gamma(t)=1-\alpha_{t}$ we obtain a loss similar to (5) without the $\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}$ weighting. Note that there remains a difference in the sampling distribution of $x_{t}$: in the masked diffusion forward process, tokens are sampled independently and do not necessarily yield exactly $(1-\alpha_{t})N$ mask tokens at time $t$, though the expected number is $(1-\alpha_{t})N$. One might be interested in whether the uniform weighting can be recovered by selecting an appropriate schedule $\alpha_{t}$. However, solving $\alpha_{t}$ such that $\alpha_{t}^{\prime}=\alpha_{t}-1$ yields $\alpha_{t}=ce^{t}+1$ and there is no $c$ that satisfies both $\alpha_{0}=1$ and $\alpha_{1}=0$. This shows that training with the MaskGIT loss (27) may not be faithfully optimizing the model likelihood.
Discrete flow matching [48].
For the linear schedule $\alpha_{t}=1-t$, our reverse transition rate matrix (8) conditioned on $x_{0}$ is:
This is the same as the conditional reverse transition rate used in Campbell et al. [48, Eq. (22)]—note that their time $t$ is reversed, and the rate matrix was therefore in the form $R^{x_{0}}(t)=\frac{1}{1-t}e_{m}(x_{0}-e_{m})^{\top}$.
SDDM [30].
Sun et al. [30] proposed a pseudo-likelihood-like objective for training discrete diffusion models that can also be applied to masked diffusion. However, their objective encounters the same challenge as Campbell et al. [29] — requiring $N$ passes of the mask prediction model. To mitigate this, they introduced a new transformer architecture, which unfortunately leads to some performance degradation.
Blackout diffusion [49].
Santos et al. [49] proposed a “blackout” diffusion process that gradually diffuses images to a black state. While this approach is similar to masked diffusion on binary data, key differences emerge when dealing with larger state spaces. In their method, image pixel intensities gradually fade out, whereas ours directly transition to a mask state. Our method offers more flexibility, being applicable to general discrete state spaces without requiring predefined structural relationships. It also demonstrates competitive performance in image generation, achieving this without knowing pixel value proximity.
SUNDAE [50, 72].
Unlike masked diffusion, SUNDAE uniformly corrupts data with random tokens in the vocab (known as uniform discrete diffusion [14]). Additionally, it uses a second loss term from cross entropy between clean data and 1-step unrolled model prediction. Similar ideas have been proposed in [73].
Appendix I Details for state-dependent rates
I.1 Derivations and time continuous limit
All derivations in this section assume that $x_{t}$ is a single token, while for $N$ tokens the masked diffusion with state-dependent rates factorises across the $N$ tokens. Learning from data of $N$ tokens using variational inference is discussed in Sec. I.2.
Given the forward transition $q(x_{t}|x_{s})$ and marginal $q(x_{s}|x_{0})$ derived in main text (Sec. 6) The reversal given $x_{0}$ is $q(x_{s}|x_{t},x_{0})=\mathrm{Cat}(x_{s};\bar{R}^{x_{0}}(t,s)^{\top}x_{t})$ for
@@ I.1 Derivations and time continuous limit
Then, combining (29) and (30) to write $q(x_{s}|x_{t}=m,x_{0})$ in an unified way yields the expression (11) in the main Sec. 6. In the second case, when $x_{t}=x_{0}\neq m$, $q(x_{s}|x_{t}\neq m,x_{0})$ from (28) simplifies dramatically and it becomes $q(x_{s}|x_{t}\neq m,x_{0})=x_{t}^{\top}x_{s}$ which is a point mass that sets $x_{s}=x_{t}$.
@@
I.2 Training and gradient estimation
The model is applied to data consisted of $N$ tokens where $x_{0}=(x_{0}^{1},\ldots,x_{0}^{(N)})$ and where each state in the masked diffusion is $x_{t}=(x_{t}^{1},\ldots,x_{t}^{(N)})$. The reverse generated model has a factorizing transition conditional of the form $\prod_{n=1}^{N}p_{\theta}(x_{s}^{(n)}|x_{t})$ where $p_{\theta}(x_{s}^{(n)}|x_{t})=q(x_{s}^{(n)}|x_{t}^{(n)},\mu_{\theta}^{(n)}(x_{t},t))$ has a form that depends on whether $x_{t}^{(n)}=m$ or $x_{t}^{(n)}\neq m$. For the first case:
@@ I.2 Training and gradient estimation
where $\mu^{(n)}_{\theta}(x_{t},t)=\text{softmax}(f_{\theta}(x_{t}))$ is a $m+1$ dimensional probability vector modelled by a NN (where the final value is constrained to be zero since $\mu^{(n)}_{\theta}(x_{t},t)$ is a reconstruction of $x_{0}^{(n)}$ which cannot be mask, so in practice the NN classifier needs to have a softmax output only over the $m$ actual token classes). Crucially, note that the NN classifier receives as input the full state $x_{t}$ of all tokens, while additional time features to encode $t$ are also included. When $x_{t}^{(n)}\neq m$ the reverse transition model is set to be $p_{\theta}(x_{s}|x_{t}^{(n)}\neq m,\{x_{t}^{(k)}\}_{k\neq n})=(x_{t}^{(n)})^{\top}x_{s}^{(n)}$ which matches precisely $q(x_{s}^{(n)}|x_{t}^{(n)}=m,x_{0}^{(n)})=(x_{t}^{(n)})^{\top}x_{s}^{(n)}$ from the forward process.
@@
Given that each $\alpha_{t,i}=1-t^{w_{i}}$, the reverse model becomes
@@
$-\frac{1}{2t}\left(\nabla_{w}\log q(x_{t}^{1}|x_{0})-\nabla_{w}\log q(x_{t}^{2}|x_{0})\right)\left[f(x_{t}^{1},x_{0})-f(x_{t}^{2},x_{0})\right].$
@@
Appendix J Experimental Details
@@
We used the continuous-time ELBO and drew one sample of $t$ for each data to estimate the integral. To reduce the variance of training, we used the same antithetic sampling trick described in Kingma et al. [33] for continuous diffusion models. We used the linear masking schedule $\alpha_{t}=1-t$ and added a small shift $\epsilon=10^{-4}$ when $t$ is close to 00 and $1$ to ensure numerical stability. The shifted schedule is $\alpha_{t}=(1-2\epsilon)(1-t)+\epsilon$. The shift leads to a support mismatch between $q(x_{1}|x_{0})$ and the prior $p(x_{1})$, leading to an undefined KL divergence term. We explain in app. E how to modify the prior distribution to allow small uniform probabilities in non-mask states to mitigate this problem. The shift leads to a non-zero reconstruction term and KL divergence term for the prior distribution but both are of negligible scale so we can safely ignore them when reporting the ELBO.
@@
J.2 OpenWebText
@@
J.3 Images
We used the same linear masking schedule as in previous experiments in all likelihood results. We used the same U-Net plus self-attention architectures from the continuous diffusion model described in Kingma et al. [33] for CIFAR-10, except that we did not use Fourier feature inputs and added an additional input embedding layer with embedding size the same as the hidden dimension of the model. For ImageNet $64\times 64$, we reduced the number of residual blocks (in one side of the U-Net structure) from 64 to 48 and added a 12-layer diffusion transformer [76] with 768 hidden dimension and 12 heads in the middle.
For both datasets we used AdamW optimizer and trained for 2M iterations. We used learning rate 0.0004, batch size 256, weight decay factor 0.01 for CIFAR-10 and learning rate 0.0002, batch size 512, weight decay factor 0.03 for ImageNet 64$\times$64. The learning rate follows a cosine annealing after 100 warm up steps. Our CIFAR-10 model is trained on 32 TPU-v5 lite for 24 hours. Our ImageNet-$64\times 64$ model is trained on 256 TPU-v5 lite for 3.5 days.
We have observed that the cosine schedule leads to better sample quality so we used it to train a cheaper model for sample visualization. This model differs from the one that achieves best likelihood in that we used 8 residual blocks (in one side of the UNet structure) and a 20-layer diffusion transformer in the middle. All other configurations are kept the same.
Appendix K Additional Results
K.1 Sample quality evaluation by GPT-2
We use the GPT-2 large model to evaluate the perplexity of samples generated by our model, following Lou et al. [32]. Results are shown in Fig. 7.
Figure 7: Generative perplexity evaluated by GPT-2 Large following Lou et al. [32]. We compare MD4 against the GPT-2 checkpoint (autoregressive baseline) and SEDD (the previous best discrete diffusion model on this task) in generating 1024-token text sequences. We investigate the effects of two orthogonal factors on sample quality: model size and decoding steps. The numbers for GPT-2 and SEDD are from Lou et al. [32].
K.2 Perplexity on OpenWebText validation set
Tab. 5 reports the final perplexity number achieved on OpenWebText validation set, corresponding to Fig. 4.
Table 5: Perplexity on OpenWebText validation set.
Size Method Perplexity ($\downarrow$)
Small Gaussian Diffusion $\leq$ 27.28
SEDD Absorb (reimpl.) $\leq$ 24.10
MD4 (Ours) $\leq$ 22.13
GenMD4 (Ours) $\leq$ 21.80
Medium MD4 (Ours) $\leq$ 16.64
K.3 Additional unconditional generation from MD4 trained on ImageNet 64$\times$64
We provide more unconditional generation results from our pixel-level modeling experiments on ImageNet 64$\times$64 in Fig. 8.
Figure 8: More unconditional samples from MD4 trained on ImageNet 64$\times$64.
K.4 Additional unconditional generation from MD4 trained on OpenWebText
Below we include two unconditioned text samples generated by our MD4 Medium model trained on OpenWebText.
K.4.1 MD4-M unconditional sample 1: 1024 tokens
@@
K.5 Conditional generation from MD4 trained on OpenWebText
We share conditionally generated text samples by MD4 Medium in Fig. 9 and observe that slow unmasking near $t=1$, enabled by the cosine schedule, tends to help produce more consist and meaningful samples than uniform unmasking counterpart.
Figure 9: Conditionally generated text samples from MD4-M. Top: MD4-M trained with linear masking schedule; Bottom: MD4-M trained with cosine masking schedule. Context text shown in blue, model-generated text in black.
K.6 Effect of discretization on zero-shot perplexity
We carried out ablation study on the effect of discretization on zero-shot perplexity. Results are included in Tab. 6. Note that this is an inference ablation with the same trained model (MD4-S trained with the continuou-time objective).
Table 6: Effect of discretization on zero-shot perplexity.
Size Timesteps LAMBADA WikiText2 PTB WikiText103 IBW
Small T = 100 $\leq$ 49.8 $\leq$ 36.1 $\leq$ 105.2 $\leq$ 36.1 $\leq$ 70.3
T = 1000 $\leq$ 48.5 $\leq$ 35.0 $\leq$ 102.5 $\leq$ 35.0 $\leq$ 68.4
T = 10000 $\leq$ 48.4 $\leq$ 34.9 $\leq$ 102.4 $\leq$ 34.9 $\leq$ 68.2
T = $\infty$ (continuous) $\leq$ 48.4 $\leq$ 34.9 $\leq$ 102.3 $\leq$ 35.9 $\leq$ 68.1